# Extreme value distribution for GRB prompt data

### v1

@@ 4 GEV analysis
The choice of the block length to extract the maxima, as discussed in Sect. 2, is not driven by first principles and is a trade-off between the goodness of the GEV distribution fit vs. the variance of the prediction. The higher the number of events per block (i.e., the longer the block duration) the better is the GEV distribution fit and the worse is the variance on the prediction. In our cases, block length of 30 days seems to offer a good compromise given the average number of events per day in both catalogues ( day). For the BATSE catalogue we obtained a sample of 66 extreme values. The richer GBM catalogue instead allowed us to derive a sample 207 extreme values. The selected ”extreme” events are shown in Fig. 3.
@@
We performed a Bayesian analysis and adopted improper uniform priors for the model parameters. The resulting posterior is proper as long as the sample size is larger than 3 (Northrop & Attalides, 2016). Sampling is carried out by the “No-U-Turn Sampler” extension (Hoffman & Gelman, 2014) to the Hamiltonian Monte Carlo algorithm (Brooks et al., 2011). The sampled chains in all cases converge to equilibrium and the posterior distribution of parameters shows a peaked distribution for the three parameters of the model. Corner plots for the six cases considered in this work are shown in Fig. 4 and summary statistics for the GEV distribution fitted parameters are reported in Table 1.
@@
In several cases the shape parameter distributions include the zero values, the simple Gümbel distribution, in their possible ranges. However, some of the distributions seem to be better fit by negative shape parameters, although in no case the evidence is very strong. This essentially implies that the available datasets are not able to unambiguously indicate whether their extreme value distributions are characterized by an asymptotic maximum value or are likely ever increasing going to an infinite number of blocks. There are also instrumental limitations to consider in evaluating the obtained results. E.g., the difficulties in recording long uninterrupted observations makes the lack of very long duration GRBs in the considered catalogues, at least partly, artificial, likely pushing the shape parameter to negative values.
@@
Since the fluence and peak flux of GRB 221009A are highly discrepant with respect to the population of GRBs from the GBM catalogue (see also Sect. 5), we also carried out the fit excluding the BOAT from the dataset. This is, at this stage, a rather arbitrary choice but, quite interestingly, without GRB 221009A the GEV distribution for the fluence would indicate an upper boundary value at erg . This figure is close to the fluence recorded for the BOAT. The analogous limit for the peak flux is instead much higher than the recorded value and, as expected, the GEV distribution for the duration is almost unaffected by the inclusion of GRB 221009A in the analysis.
@@
As discussed in Sect. 2, the fit quality was checked by diagnostic plots, as the widely known QQ plots. For the both the datasets considered in this work, the distributions of durations, peak fluxes and fluences turn out to be nicely fit by a GEV distribution (Fig. 1), with the already mentioned notable exception of GRB 220109A for the fluence and the peak flux that stands clearly out among the extreme values from the GBM catalogue. As already pointed out in past analyses (e.g., Burns et al., 2023; Malesani et al., 2023), the peculiarity of GRB 220109A is due to a combination of factors, i.e., persistent high flux emission, long duration and the spectral hardness.
@@
Finally, the return plots are shown in Fig.2. The reported levels are expected to be exceeded on average once every years, i.e., the return level is exceeded in any particular block with probability .
@@
For the BATSE catalogue, as expected from the analysis of the QQ plot, the return levels are all well consistent with the modelling (Fig. 2). Things are considerably different for the GBM catalogue data, where again GRB 220109A appears grossly inconsistent with the model.
@@
Just extrapolating the fit distribution, the median value of the model return level, at the fluence of GRB 220109A, is more than years, and the BOAT fluence enters the 95% allowed range at years. A peak flux value comparable to what shown by GRB 220109A is expected in years, while the observed peak flux enters the 95% range in years. Removing GRB 220109A the GEV distributions for the fluence and peak flux give a negligible probability of observing an event of comparable fluence or peak flux for time scales of thousands years.
@@ 5 Discussion and Conclusions
The most interesting feature of a GEV analysis is that it does not require any assumptions beyond that the analysed data are drawn from a stationary distribution. Durations, peak fluxes and fluences for the brightest events detected by a specific instrument are satisfying this requirements, since, to a large extent, we can assume that the natural degradation of the instruments have a minor or even negligible impact on their capability to detect observationally bright events. In addition, focusing on the longest, brightest and with the highest fluence events, we are also automatically excluding from the data sets GRBs belonging to the short duration class (Kouveliotou et al., 1993) with the partial exception of the extreme values for the peak fluxes. Quite interestingly, for both the considered catalogues, a percentage of 14% (BATSE) and (GBM) of peak flux extreme values is due to events shorter than the canonical 2 s value. On the other hand, in spite of the fact that the brightest events are also, unavoidably, the best studied, there is still the (more or less virtual) possibility that the BATSE and GBM samples are polluted by very bright events due to phenomena erroneously classified as GRBs. As a matter of fact, it was also proposed that some of the brightest events come from a different (and subdominant) population of GRBs (e.g., Finke & Razzaque, 2024; Briel et al., 2025). Analogous considerations have been proposed for the ultra-long GRBs ( 1000 s) (e.g., Levan et al., 2014) that, however, are not included in the present analysis.
@@
Our results indicate that, apart from the case of GRB 221009A, that we are going to discuss later, the GEV distributions for GRB reported in the BATSE and GBM catalogues are consistent with being drawn from a single population of events. Their respective GEV distributions are also substantially consistent with each other (Table 1). For instance, the return level for a GRB duration of s is close to one event in 20-30 years in both cases, possibly inconsistent with the few ultra-long GRBs observed so far, although the inefficient recording of very long duration events makes any conclusions about this topic at this stage highly preliminary.
@@
As already observed, the data for GRB 221009A are well consistent with the population GEV distributions if we refer to the duration. They are instead highly discrepant for the fluence and peak flux. This is a non trivial point, and shows that the peculiarities of the BOAT are due to the combination of several factors as, long-lasting high-flux emission (and its spectral hardness, e.g., Frederiks et al., 2023) rather than to a very long or very bright flux only. This is also in agreement with findings from previous works (e.g., Malesani et al., 2023; Burns et al., 2023). The conclusions about the peak flux hold even if different baselines for the flux computations are used.
@@
The fluence and peak flux of GRB 221009A are, therefore, grossly inconsistent with the GEV distribution. The median of the return level for the fluence yields a probability of having one event of this fluence in about 1000 years, while for the peak flux the median return value is about 140 years. All in all, GRB 221009A appears an unlikely event for its fluence and peak flux, given the duration of the monitoring of the considered missions. If we, on the contrary, adopt the hypothesis that an event with fluence and peak flux as large as or larger than GRB 221009A is unlikely, this might imply that one of the assumptions of the present analysis, i.e., that the measured extreme values are drawn from an homogenous population of GRBs, is violated. As a matter of fact, the possibility that GRB 221009A is not a typical cosmological GRB has been proposed and discussed in the literature. This might be due to a different geometry of the event (Finke & Razzaque, 2024), a misclassification of the event as a GRB (Navia et al., 2024) or just low-probability intense gravitational lens enhancement (Bloom, 2022). In all cases, these hypotheses require that their rate of occurrence is low enough not to affect the bulk of the GRB population observations and that they affect mainly the fluence, and to a lower extent the peak flux, but not duration of the detected events.
@@
An analysis considering all the GRBs detected by any instrument after a proper homogenization is in principle possible, although the possibility of introducing artefacts due to the incomplete knowledge of emission spectra and the intrinsic limitations of the involved instruments is high. We leave this possible development to a future work.
@@ 6 Data availability
We developed software tools and used third-party libraries coded in the julia language (v. 1.13.4-1.13.5, Bezanson et al., 2017)444https://julialang.org/. GEV models and fitting tools are from the Extremes package555https://jojal5.github.io/Extremes.jl/dev/ (v. 1.0.3-1.0.5, Jalbert et al., 2024). Plots are produced by the makie666https://docs.makie.org/v0.22/ (v. 0.22.2-0.24.0, Danisch & Krumbiegel, 2021) with the CairoMakie backend, and corner plot are produces by the PairPlots777https://sefffal.github.io/PairPlots.jl/dev/ (v. 3.0.1, Thompson, 2023) packages. Project management was carried out by the DrWatson888https://juliadynamics.github.io/DrWatson.jl/dev/ package (v. 2.18.0, Datseris et al., 2020). Bibliographic research has also been aided by the Google Gemini (v. 2.5) IA model.999https://gemini.google.com/

### v2

@@ 4 GEV analysis
The choice of the block length to extract the maxima, as discussed in Sect. 2, is not driven by first principles and is a trade-off between the goodness of the GEV distribution fit versus the variance of the prediction. The higher the number of events per block (i.e. the longer the block duration), the better the GEV distribution fit and the worse is the variance on the prediction. In our cases, ablock length of 30 days seems to offer a good compromise given the average number of events per day in both catalogues ( day). For the BATSE catalogue, we obtained a sample of 66 extreme values. The richer GBM catalogue instead allowed us to derive a sample of 207 extreme values. The selected ’extreme’ events are shown in Fig. 3.
@@
We performed a Bayesian analysis and adopted improper uniform priors for the model parameters. The resulting posterior is proper as long as the sample size is larger than 3 (Northrop & Attalides, 2016). Sampling is carried out by the ‘No-U-Turn Sampler’ extension (Hoffman & Gelman, 2014) to the Hamiltonian Monte Carlo algorithm (Brooks et al., 2011). The sampled chains in all cases converge to equilibrium, and the posterior distribution of parameters shows a peaked distribution for the three parameters of the model. Corner plots for the six cases considered in this work are shown in Fig. 4, and summary statistics for the GEV-distribution fitted parameters are reported in Table 1.
@@
In several cases, the shape parameter distributions include the zero values and the simple Gümbel distribution in their possible ranges. However, some of the distributions seem to be better fitted by negative shape parameters, although in no case is the evidence very strong. This essentially implies that the available data sets are not able to unambiguously indicate whether their extreme value distributions are characterised by an asymptotic maximum value or are likely ever increasing going to an infinite number of blocks. There are also instrumental limitations to consider in evaluating the obtained results. For example, the difficulties in recording long uninterrupted observations makes the lack of very long duration GRBs in the considered catalogues, at least partly, artificial, likely pushing the shape parameter to negative values.
@@
Since the fluence and peak flux of GRB 221009A are highly discrepant with respect to the population of GRBs from the GBM catalogue (see also Sect. 5), we also carried out the fit excluding the BOAT from the data set. This is, at this stage, a rather arbitrary choice, but, quite interestingly, without GRB 221009A the GEV distribution for the fluence would indicate an upper boundary value at erg cm-2. This figure is close to the fluence recorded for the BOAT. The analogous limit for the peak flux is instead much higher than the recorded value, and, as expected, the GEV distribution for the duration is almost unaffected by the inclusion of GRB 221009A in the analysis.
@@
As discussed in Sect. 2, the fit quality was checked by diagnostic plots, as were the widely known QQ plots. For both the data sets considered in this work, the distributions of durations, peak fluxes, and fluences turn out to be nicely fit by a GEV distribution (Fig. 1), with the already-mentioned notable exception of GRB 220109A for the fluence and the peak flux that clearly stands out among the extreme values from the GBM catalogue. As already pointed out in past analyses (e.g. Burns et al., 2023; Malesani et al., 2023), the peculiarity of GRB 220109A is due to a combination of factors; for example, persistent high flux emission, long duration, and the spectral hardness.
@@
Finally, the return plots are shown in Fig.2. The reported levels are expected to be exceeded on average once every years; i.e., the return level is exceeded in any particular block with a probability of .
@@
For the BATSE catalogue, as expected from the analysis of the QQ plot, the return levels are all well consistent with the modelling (Fig. 2). Things are considerably different for the GBM catalogue data, where GRB 220109A again appears grossly inconsistent with the model.
@@
Just extrapolating the fit distribution, the median value of the model return level, at the fluence of GRB 220109A, is more than years, and the BOAT fluence enters the 95% allowed range at years. A peak flux value comparable to what was shown by GRB 220109A is expected in years, while the observed peak flux enters the 95% range in years. Removing GRB 220109A the GEV distributions for the fluence and peak flux gives a negligible probability of observing an event of comparable fluence or peak flux for timescales of thousands years.
@@ 5 Discussion and conclusions
The most interesting feature of a GEV analysis is that it does not require any assumptions beyond that the analysed data are drawn from a stationary distribution. Durations, peak fluxes, and fluences for the brightest events detected by a specific instrument satisfy this requirement, since, to a large extent, we can assume that the natural degradation of the instruments have a minor or even negligible impact on their capability to detect observationally bright events. In addition, focusing on the longest, brightest, and highest fluence events, we are also automatically excluding GRBs belonging to the short-duration class from the data sets (Kouveliotou et al., 1993), with the partial exception of the extreme values for the peak fluxes. Quite interestingly, for both the considered catalogues, 14% (BATSE) and (GBM) of peak flux extreme values are due to events shorter than the canonical 2 s value. On the other hand, in spite of the fact that the brightest events are also, unavoidably, the best studied, there is still the (more or less virtual) possibility that the BATSE and GBM samples are polluted by very bright events due to phenomena erroneously classified as GRBs. As a matter of fact, it was also proposed that some of the brightest events come from a different (and sub-dominant) population of GRBs (e.g. Finke & Razzaque, 2024; Briel et al., 2025). Analogous considerations have been proposed for the ultra-long GRBs ( 1000 s) (e.g. Levan et al., 2014), which, however, are not included in the present analysis.
@@
Our results indicate that, apart from the case of GRB 221009A, which we discuss in the next paragraph, the GEV distributions for the GRB reported in the BATSE and GBM catalogues are consistent with being drawn from a single population of events. Their respective GEV distributions are also substantially consistent with each other (Table 1). For instance, the return level for a GRB duration of s is close to one event in 20-30 years in both cases, which is possibly inconsistent with the few ultra-long GRBs observed so far, although the inefficient recording of very-long-duration events makes any conclusions about this topic at this stage simply preliminary.
@@
As already observed, the data for GRB 221009A are well consistent with the population GEV distributions if we refer to the duration. They are instead highly discrepant for the fluence and peak flux. This is a non-trivial point, and it shows that the peculiarities of the BOAT are due to the combination of several factors such as long-lasting high-flux emission (and its spectral hardness, e.g. Frederiks et al., 2023) rather than to a very long or very bright flux only. This is also in agreement with findings from previous works (e.g. Malesani et al., 2023; Burns et al., 2023). The conclusions about the peak flux hold even if different baselines for the flux computations are used.
@@
The fluence and peak flux of GRB 221009A are, therefore, grossly inconsistent with the GEV distribution. The median of the return level for the fluence yields a probability of one event of this fluence occurring in about 1000 years, while for the peak flux the median return value is about 140 years. All in all, GRB 221009A appears an unlikely event for its fluence and peak flux, given the duration of the monitoring of the considered missions. On the contrary, if we adopt the hypothesis that an event with fluence and peak flux as large as or larger than GRB 221009A is unlikely, this might imply that one of the assumptions of the present analysis –i.e. that the measured extreme values are drawn from a homogenous population of GRBs– is violated. As a matter of fact, the possibility that GRB 221009A is not a typical cosmological GRB has been proposed and discussed in the literature. This might be due to a different geometry of the event (Finke & Razzaque, 2024), a misclassification of the event as a GRB (Navia et al., 2024), or just low-probability intense gravitational-lens enhancement (Bloom, 2022). In all cases, these hypotheses require their rate of occurrence to be low enough not to affect the bulk of the GRB population observations and to mainly affect the fluence and, to a lesser extent the peak flux, but not the duration of the detected events.
@@
An analysis considering all the GRBs detected by any instrument after a proper homogenisation is in principle possible, although the possibility of introducing artefacts due to incomplete knowledge of emission spectra and the intrinsic limitations of the involved instruments is high. We leave this possible development to a future work.
@@ 6 Data availability
We developed software tools and used third-party libraries coded in the julia language (v. 1.13.4-1.13.5, Bezanson et al., 2017).444https://julialang.org/ GEV models and fitting tools are from the Extremes package555https://jojal5.github.io/Extremes.jl/dev/ (v. 1.0.3-1.0.5, Jalbert et al., 2024). Plots are produced by the makie666https://docs.makie.org/v0.22/ (v. 0.22.2-0.24.0, Danisch & Krumbiegel, 2021) with the CairoMakie backend, and corner plot are produces by the PairPlots777https://sefffal.github.io/PairPlots.jl/dev/ (v. 3.0.1, Thompson, 2023) packages. Project management was carried out by the DrWatson888https://juliadynamics.github.io/DrWatson.jl/dev/ package (v. 2.18.0, Datseris et al., 2020). Bibliographic research has also been aided by the Google Gemini (v. 2.5) IA model.999https://gemini.google.com/